The odds against a husband who is 45 years old, living till he is 70 are
7:5 and the odds against his wife who is now 36, living till she is 61
are 5:3 . Find the probability that a) the couple will alive 25 years hence
b) exactly one of them will be alive 25 years hence c) none of them will
be alive 25 years hence d) at least one of them will be alive 25 years
hence

Grade:12

2 Answers

jitender lakhanpal
10 years ago

P(H)=5/12    -- the probab that husband lives 25 years hence

P(W)=3/8    -- the probab that wife lives 25 years hence

a) the couple lives 25 years hence is P(H)*P(W) = 5/32

b) exactly one of them lives 25 years hence = P(H) * P(Wbar) + P (W) * P(Hbar)

c)none of them be alive P(Hbar) *P(Wbar)

d) atleast one of them alive 1-(probab none of them alive)

We are  here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

Thanks and Regards

Jitender

(askIITians expert)

Jagmal Singh
5 years ago
 

P(H)=5/12    -- the probab that husband lives 25 years hence

P(W)=3/8    -- the probab that wife lives 25 years hence

a) the couple lives 25 years hence is P(H)*P(W) = 15/96

b) exactly one of them lives 25 years hence = P(H) * P(Wbar) + P (W) * P(Hbar) = 46/96

c)none of them be alive P(Hbar) *P(Wbar) = 35/96

d) atleast one of them alive 1-(probab none of them alive) = 1-(35/96) = 61/96